Shantelsurry:

If QS and TV are parallel lines and mSRU = 50°, what is mTUR?

bruhimagirl22222:

maybe try solving this and tell me if u understand

1 attachment
shayy:

I don’t understand

xXQuintonXx:

i cant tell if its asking if those 2 r parrarel lines or not

Bonnnnn:

180-51 a straight line == 180 and you already have 51

shayy:

It’s 5

shayy:

50

bruhimagirl22222:

ok and how did u get 50?

shayy:

That’s the measurements

bruhimagirl22222:

do u have a diagram

Alexxxxxx:

what are you trying to find for this?????

jhonyy9:

on this posted image point U dont exist please check these posted details

jhonyy9:

ohhh who posted it is offline

jhonyy9:

If QS and TV are parallel lines and mSRU = 50°, what is mTUR? looking on these details and the posted image i think that this problem wan be in this way given m<POR = 51° and the question find m<TRQ = ?

jhonyy9:

any idea ?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 m<POR = 51° => m<ORQ = 51° => m<TRS = 51° m<TRQ +m<ORS +m<ORQ +m<TRS = 360° m<TRQ =m<ORS m<ORQ=m<TRS m<TRQ +m<ORS + 51 +51 = 360 x +x = 360 -102 2x = 258 x = 258/2 x = 129 m<TRQ = 129° hope helped understandably easy \(\color{#0cbb34}{\text{End of Quote}}\)